# Quantifying Uncertainty in Error Consistency: Towards Reliable Behavioral Comparison of Classifiers

### v1

@@ 3.2 Using Proposition 2 for simulations
Proposition 2 provides such a model in which the underlying marginals and copy probability can be specified to generate data. Note that, depending on the copy probability, the observed marginals will deviate from the underlying marginal: The more responses the second classifier copies from the first, the more its marginal distribution adapts. To simulate data with specific marginal distributions $P^{(1)}$ and $P^{(2)}$ as well as EC, one can simulate data with copy probability and underlying marginals
@@
Equipped with this generative model, we can estimate confidence intervals for a pair of classifiers as follows: We first sample $N$ trials for the first classifier by drawing from a binomial distribution $r^{(1)}_{i}\sim\tilde{P}^{(1)}$. Next, we copy $p_{copy}\cdot N$ responses from the first classifier to the second. For the remaining $(1-p_{copy})\cdot N$ responses, we sample independently from a binomial distribution from the underlying marginal $r^{(2)}_{i}\sim\tilde{P}^{(2)}$. This yields one data set based on which we estimate EC. Repeating this process then yields an estimate for the variability of these EC values.
We plot the resulting confidence intervals in Figure 3, and include sample sizes used in the literature (Geirhos et al., 2021; Ollikka et al., 2024; Li et al., 2025; Wiles et al., 2024) as reference points. Note that the x-axis is on a logarithmic scale and reaches values that are very difficult to meet in experiments with human observers. Even so, the width of the confidence intervals (especially that of classifiers close to floor / ceiling performance) is very large and shrinks only slowly, covering $0.2$ units of $\kappa$ even at $N=400$ trials, which is roughly the size of the difference found between humans and machines by Geirhos et al. (2021). Especially noteworthy is the fact that at lower $N$ and performance close to floor / ceiling, the bootstrapped ECs are biased and undershoot the true value. This happens for a simple reason: If one classifier is always correct (or always wrong) the EC will be zero irrespective of the accuracy of the other one. In the critical performance regimes and at a low number of trials, sub-sampling trial sequences for which these edge cases occur becomes likely enough to bias results.
@@
A core limitation of our work is that we do not present an analytic solution for confidence intervals, but rely on bootstrapping instead. This also limits the numerical accuracy of the significance tests, which cannot resolve p-value differences $\leq\frac{1}{N}$. While the analytic solutions (along the lines of Gwet (2016), Fleiss et al. (1969), and Vanbelle et al. (2024)) could improve this slightly, we believe that for the discussed purpose here, our methodology is the pragmatic choice that comes with less restrictions (and certainly much better than not reporting confidence intervals at all).
@@ 3.3 Significance Tests for Error Consistency
Now that we can quantify the degree of uncertainty, we can also derive significance tests for error consistency. Significance tests are an invaluable tool for empirical scientists because they provide a standardized way of checking if results are statistically reliable. Given two sequences of binary trials of length $N$ which lead to an EC of $\kappa$, we want to calculate a p-value for $\kappa$. The p-value is the probability of observing an event at least as extreme as the observed one under the null hypothesis of independent binomial observers.
To obtain the distribution of EC values obtained under the null hypothesis, we first need to model the independent binomial observers. They are fully characterized by their individual accuracies, of which the empirically observed accuracies provide an estimate. The reliability of this estimate depends on the number of trials. If we assume uniform priors over these accuracies, the posteriors are beta-distributions characterized by the number of (in)correctly solved trials, $\mathrm{Beta}(\sum_{i}r^{(j)}_{i},N-\sum_{i}r^{(j)}_{i})$.
We can now sample the two individual accuracies from their respective posteriors, sample $N$ trials from independent binomial distributions and calculate the resulting EC. Repeating this procedure $M$ times leads to a distribution of EC values expected under the null hypothesis. From this distribution, we then calculate how many times the absolute value of the simulated EC exceeded the absolute value of the empirically measured EC, for a two-sided significance test. We demonstrate this method by applying it to data from model-vs-human in Appendix F to obtain p-values of the error consistency between a human subject and the best-performing model, a CLIP-trained ViT (Radford et al., 2021).
@@ 4.1 Model-VS-Human
Next, we re-analyze well-known literature results by calculating confidence intervals around their empirically measured values. At the focus of our attention is the work by Geirhos et al. (2021), where a cohort of human observers was asked to classify corrupted natural images. The same images were also shown to 52 DNNs, to evaluate their alignment to human behavior. EC values were calculated per condition and averaged as described in Section 3. From our analysis in Figure 3, we already know that some of the error consistencies were calculated in a low-data regime that will result in large confidence intervals around the individual EC values, but we now compute confidence intervals specifically for their empirical data. We do this by bootstrapping through the entire calculation, to arrive at an empirical distribution of final scores, which we plot for humans and every model in Figure 4. The best model is a CLIP-trained ViT(Radford et al., 2021) which performs much better on the distribution shifts like sketches and stylized images. For each bootstrap, we also obtain a ranking of models. Given that the CIs are fairly wide (for example, no model’s confidence interval is separated from that of its neighbors), we wonder whether different bootstraps imply different model rankings. To check this, we calculate Kendall’s $\tau$ between the original model ranking and the ranking implied by each of the 10k bootstraps. $\tau$ is a statistical measure of the rank correlation between two ordered sequences, similar to Spearman’s rank correlation, but with the benefit of an intuitive interpretation: $\tau$ measures the difference between the probability that a pair of observations is concordant and the probability that it is discordant. The average $\tau$ is $0.88$, which means that two randomly sampled pairs have an 88% chance of being concordant. Evidently, the main conclusion drawn by Geirhos et al. (2021) clearly still holds: Humans are more consistent among themselves than models are to humans. Nevertheless, adding confidence intervals to the reported EC values better quantifies the degree of uncertainty about the measurements for individual models, and the 95% CIs do indeed overlap, meaning that differences between models are not resolved.
Figure 4: Models from MvH ranked by their EC to humans. Inter-human EC is plotted in red. We plot the 95% percentile interval for the mean EC obtained via bootstrapping, using the standard exclusion criteria of the benchmark (orange and black points in Figure 2b). For the mapping of model indices to names, see Table 2.
@@ 4.2 Brain-Score
One of the most widely-known applications of EC is the influential Brain-Score benchmark [^1] https://www.brain-score.org (Schrimpf et al., 2018), which evaluates arbitrary DNNs by how much they resemble the brain. One dimension of this evaluation is the behavioral similarity to humans, including error consistency as measured via the data by Geirhos et al. (2021). Here we assess how reliable the ranking of models implied by their EC to humans really is.
In Brain-Score, the EC for each condition is first divided by the human-to-human EC for this condition, to express values relative to a ceiling. We begin by reverting this operation to obtain raw EC values. To properly bootstrap EC values for these models, we would need access to the complete sequences of trials generated by the models, but this data is not stored anywhere, and re-running the models (not all of which are publicly available anyway) seems excessive. Hence, we instead make mild assumptions to estimate conservative CIs around the empirical values, demonstrating the utility of our copy-model. (See Appendix E for details on this procedure.) We obtain the result depicted in Figure 5 for the top-30 models (by EC to humans). Again, the best model is a CLIP-trained ViT. For the other models, the CIs are so large that any of the top ten models could be the second-best one, meaning that the true rank order between models remains unclear.
@@
Figure 5: Confidence Intervals for Brain-Score ECs. Like in Figure 4, we plot the 95% confidence interval of the EC of the top 30 Brain-Score models.
@@ 5 Discussion
In this work, we have built on the error consistency metric proposed by Geirhos et al. (2020), demonstrating how confidence intervals and significance tests for empirical measurements of EC can be calculated using bootstrapping techniques. Additionally, we have presented a new computational model of EC for planning sufficiently powerful, conclusive experiments. We have used our methodology to revisit influential Neuro-AI results Schrimpf et al. (2018), finding that many results are not statistically reliable. We recommend the following practices for future behavioral benchmarking to avoid these issues using our provided python package.
5.1 Practical Recommendations: How to reliably benchmark behavioral alignment.
@@
From our analysis, we can derive concrete recommendations for practitioners. As shown in Figure 3, a sufficiently high number of trials is required to obtain small confidence intervals around EC values. As a rule of thumb, we suggest to collect at least 1000 trials per classifier to balance this requirement against practical constraints. Since the size of CIs also depends on the accuracy mismatch between classifiers, we suggest to aim for an accuracy around 75%, but not more than 90% to avoid ceiling-effects, which can be severe. The same holds for floor-effects. This is in agreement with standard best practices from psychophysics, where one aims for a similar accuracy level in human observers to sustain motivation whilst avoiding ceiling effects. Finally, future work should always report confidence intervals and check results for statistical significance, which can easily be done with the python package we provide.
5.2 Conclusion.
@@
The field of ML research relies heavily on the powerful machinery of benchmarking (Hardt, 2025). Hence, the correctness and reliability of benchmarks is crucial for progress, since ill-calibrated benchmarks will lead us astray. The field is also notorious for not quantifying uncertainty properly (Pineau et al., 2021; Miller, 2024; Lehmann and Paromau, 2025; Bouthillier et al., 2021), which is especially important in the context of benchmarks because of their large influence on other developments. Our work raises general questions about how benchmarks should aggregate results. For benchmarks that explicitly compute rankings, we give the concrete recommendations of (a) providing confidence intervals around individual estimates and (b) quantifying the stability of the resulting rankings as we did with error consistency.
@@
In this work, we have revisited results by Geirhos et al. (2021) which influence the Brain-Score benchmark Schrimpf et al. (2018), revealing that many differences reported between deep vision models are not statistically reliable. The error consistency metric is particularly susceptible to such issues, because it is inherently noisy, so the requirements for stable measurements are hard to meet in practice. We envision that our work will help the field to overcome these issues, by providing tools to plan sufficiently powerful experiments.
@@
Funded, in part, by the Collaborative Research Centre (CRC) “Robust Vision – Inference Principles and Neural Mechanisms” of the German Research Foundation (DFG; SFB 1233), project number 276693517. This work was additionally supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. WB acknowledges financial support via an Emmy Noether Grant funded by the German Research Foundation (DFG) under grant no. BR 6382/1-1, via the Open Philanthropy Foundation funded by the Good Ventures Foundation and the German Research Foundation (DFG) through the SFB 1233, Robust Vision: Inference Principles and Neural Mechanisms, TP A2, project number: 276693517. WB, KM and FAW are members of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. KM is supported by the Carl Zeiss Foundation through the CZS Institute for Artificial Intelligence and Law. This research utilized compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Thomas Klein.
We would like to thank Robert Geirhos for helpful discussions and for providing us with his raw data for model-vs-human. We would also like to thank Katharina Deckenbach for initial discussions in this project.
@@ Appendix A Related Work
Within the last decade, Deep Neural Networks (DNNs) have been proposed as models of the human visual system [Doerig et al., 2023, Kriegeskorte, 2015, Cichy and Kaiser, 2019, Kietzmann et al., 2017], not only because they form the only class of image-computable models that achieves human-level performance on benchmark tasks like ImageNet [Russakovsky et al., 2015], but also because performance-optimized DNNs are the best predictors of biological neural activations [Yamins et al., 2014, Kubilius et al., 2019, Zhuang et al., 2021]. However, the suitability of DNNs as models of vision is debated [Bowers et al., 2022, Wichmann and Geirhos, 2023, Maniquet et al., 2024, Serre, 2019], as is the question of how comparisons should best be conducted [Firestone, 2020, Lonnqvist et al., 2021, Wichmann and Geirhos, 2023].
@@
The underlying idea motivating error consistency is that of “molecular psychophysics” [Green, 1963], which is to go beyond aggregate measures like accuracy, and instead compare behavior on a trial-by-trial basis. EC promises to achieve this by building on Cohen’s $\kappa$ [Cohen, 1960, Martín Andrés and Álvarez Hernández, 2024], which considers the trial-level agreement observed for a pair of classifiers. Geirhos et al. [2020] proposed EC in the context of human-machine comparison, but in principle, the method applies to the general setting of comparing arbitrary classifiers. Martín Andrés and Álvarez Hernández [2024] observe that $\kappa$ is a biased estimator and should be corrected, which we support in our implementation. Safak [2020] also observe the problem of $\kappa^{\prime}s$ dependency on the marginals, which we outline in Section 2. They propose to scale $\kappa$ between its theoretical limits in an attempt to obtain a measure of error consistency that is orthogonal to accuracy. Goel et al. [2025] propose to apply Cohen’s $\kappa$ directly to the responses given by both classifiers $\hat{y}_{i}$ rather than the correctness values $r_{i}$, and adapt the marginal probabilities of independent classifiers to reflect their accuracy. We still focus on the standard error consistency proposed by Geirhos et al. [2020] because it has seen the widest application. In principle, our bootstrapping methodology also works for these other metrics.
@@
Benchmarking has a long history and is central to modern machine learning [Koch and Peterson, 2024, Singh et al., 2025, Hardt, 2025, Orr and Kang, 2024]. For example, the popular Papers with Code database lists approximately 14,000 different benchmarks [Code, 2025]. With the emergence of modern benchmarks and leaderboards, issues such as overfitting to benchmarks [Singh et al., 2025], the limitations of metrics [Thomas and Uminsky, 2020], and various sources of uncertainty have been identified and challenged [Lehmann and Paromau, 2025, Bouthillier et al., 2021]. In general, it has been criticized that benchmarks should not be the primary goal of scientific machine learning research [Alzahrani et al., 2024]. At the same time, it has been pointed out that benchmarking itself requires a scientific approach and some kind of standard methodology [Thiyagalingam et al., 2022, Sculley et al., 2018, Hardt, 2025]. Similar to our work, Nado et al. [2021] argued that uncertainty is an important aspect to consider for benchmarks. In contrast to previous work, we focus on uncertainty in the behavioral comparison of classifiers using error consistency.
Appendix B Proof of Proposition 2

### v2

@@ 3.2 Using Proposition˜2 for simulations
Proposition˜2 provides such a model in which the underlying marginals and copy probability can be specified to generate data. Note that, depending on the copy probability, the observed marginals will deviate from the underlying marginal: The more responses the second classifier copies from the first, the more its marginal distribution adapts. To simulate data with specific marginal distributions $P^{(1)}$ and $P^{(2)}$ as well as EC, one can simulate data with copy probability and underlying marginals
@@
Equipped with this generative model, we can estimate confidence intervals for a pair of classifiers as follows: We first sample $N$ trials for the first classifier by drawing from a Bernoulli distribution $r^{(1)}_{i}\sim\tilde{P}^{(1)}$. Next, we copy $p_{copy}\cdot N$ responses from the first classifier to the second. For the remaining $(1-p_{copy})\cdot N$ responses, we sample independently from a Bernoulli distribution from the underlying marginal $r^{(2)}_{i}\sim\tilde{P}^{(2)}$. This yields one data set based on which we estimate EC. Repeating this process then yields an estimate for the variability of these EC values. We provide an algorithmic description of this approach in Algorithm˜2.
We plot the resulting confidence intervals in Figure˜3, and include sample sizes used in the literature (geirhos2021partial; ollikka2024humans; li2025laion; wiles2024revisiting) as reference points. We focus on the most severe case where $\kappa=0.5$, but discuss the effect of $\kappa$ on CI size in Appendix˜E. Note that the x-axis is on a logarithmic scale and reaches values that are very difficult to meet in experiments with human observers. Even so, the width of the confidence intervals (especially that of classifiers close to floor / ceiling performance) is very large and shrinks only slowly, covering $0.2$ units of $\kappa$ even at $N=400$ trials, which is roughly the size of the difference found between humans and machines by geirhos2021partial. Especially noteworthy is the fact that at lower $N$ and performance close to floor / ceiling, the bootstrapped ECs are biased and undershoot the true value. This happens for a simple reason: If one classifier is always correct (or always wrong) the EC will be zero irrespective of the accuracy of the other one. In the critical performance regimes and at a low number of trials, sub-sampling trial sequences for which these edge cases occur becomes likely enough to bias results.
@@
A core limitation of our work is that we do not present an analytic solution for confidence intervals, but rely on bootstrapping instead. This also limits the numerical accuracy of the significance tests, which cannot resolve p-value differences $\leq\frac{1}{N}$. While the analytic solutions (along the lines of Gwet_2016, Fleiss_1969, and Vanbelle_2024) could improve this slightly, we believe that for the discussed purpose here, our methodology is the pragmatic choice that comes with fewer restrictions (and is certainly much better than not reporting confidence intervals at all).
@@ 3.3 Significance Tests for Error Consistency
Now that we can quantify the degree of uncertainty, we can also derive significance tests for error consistency. Significance tests are an invaluable tool for empirical scientists because they provide a standardized way of checking if results are statistically reliable. Given two sequences of binary trials of length $N$ which lead to an EC of $\kappa$, we want to calculate a p-value for $\kappa$. The p-value is the probability of observing an event at least as extreme as the observed one under the null hypothesis of two independent binomial observers, i.e. $\kappa=0$.
Basic Significance Test.
To obtain the distribution of EC values obtained under the null hypothesis $H_{0}:\kappa=0$, we first need to model the independent binomial observers that constitute the null hypothesis. They are fully characterized by their individual accuracies, of which the empirically observed accuracies provide an estimate. The reliability of this estimate depends on the number of trials. If we assume uniform priors over these accuracies, the posteriors are beta-distributions characterized by the number of (in)correctly solved trials, $\mathrm{Beta}(N-\sum_{i}r^{(j)}_{i},\sum_{i}r^{(j)}_{i})$.
We can now sample the two individual accuracies from their respective posteriors, sample $N$ trials from independent binomial distributions and calculate the resulting EC. Repeating this procedure $M$ times leads to a distribution of EC values expected under the null hypothesis. From this distribution, we then calculate how many times the absolute value of the simulated EC exceeded the absolute value of the empirically measured EC, for a two-sided significance test. We demonstrate this method by applying it to data from model-vs-human in Appendix˜H to obtain p-values of the error consistency between a human subject and the best-performing model, a CLIP-trained ViT (radford2021learning).
Advanced Significance Test.
The empirical scenario which we investigate in Section˜4 demands a slightly different significance test: Given a fixed reference observer (e.g. a human) and two candidate observers (e.g. two DNNs), is the difference of ECs to the reference observer found between the two candidates statistically significant?
Such a test is also possible. Let $\kappa_{c_{1}},\kappa_{c_{2}}$ be the error consistencies of the two candidates. We consider the null hypothesis of the candidates having the same error consistency to the reference classifier, $H_{0}:\kappa_{c_{1}}=\kappa_{c_{2}}$. We first obtain a distribution of error consistency differences via bootstrapping as above: First, we randomly sample $N$ trial indices $i$ with replacement and uniform probability. We select the corresponding response-correctness-triplets $(r^{(ref)}_{i},r^{(c_{1})}_{i},r^{(c_{2})}_{i})\in\{0,1\}^{3}$ and re-calculate EC values and their difference. To obtain the distribution of differences one would expect under $H_{0}$, we simply exchange $r^{(c_{1})}_{i},r^{(c_{2})}_{i}$ with probability $p=0.5$, as if there had been only one underlying model. We can then perform a t-test between the distributions of differences to compute p-values.
@@ 4.1 Model-vs-Human
[…]
For each bootstrap, we also obtain a ranking of models. Given that the CIs are fairly wide (for example, no model’s confidence interval is separated from that of its neighbors), we wonder whether different bootstraps imply different model rankings. To check this, we calculate Kendall’s $\tau$ between the original model ranking and the ranking implied by each of the 10k bootstraps. $\tau$ is a statistical measure of the rank correlation between two ordered sequences, similar to Spearman’s rank correlation, but with the benefit of an intuitive interpretation: $\tau$ measures the difference between the probability that a pair of observations is concordant and the probability that it is discordant. The average $\tau$ is $0.88$, which means that two randomly sampled pairs have an 88% chance of being concordant. Evidently, the main conclusion drawn by geirhos2021partial clearly still holds: Humans are more consistent among themselves than models are to humans. Nevertheless, adding confidence intervals to the reported EC values better quantifies the degree of uncertainty about the measurements for individual models, and the 95% CIs do indeed overlap, meaning that differences between models are not resolved.
Figure 4: Models from MvH ranked by their EC to humans. Inter-human EC is plotted in red. We plot the 95% percentile interval for the mean EC obtained via bootstrapping, using the standard exclusion criteria of the benchmark (orange and black points in Figure˜2b). For the mapping of model indices to names, see Table˜2. Note that while the gap to humans is large, confidence intervals of many models overlap.
@@ 4.2 Brain-Score
One of the most widely-known applications of EC is the influential Brain-Score benchmark [^1] https://www.brain-score.org (schrimpf2018brainscore), which evaluates arbitrary DNNs by how much they resemble the brain. One dimension of this evaluation is the behavioral similarity to humans, including error consistency as measured via the data by geirhos2021partial. Here we assess how reliable the ranking of models implied by their EC to humans really is.
In Brain-Score, the EC for each condition is first divided by the human-to-human EC for this condition, to express values relative to a ceiling. We begin by reverting this operation to obtain raw EC values. To properly bootstrap EC values for these models, we would need access to the complete sequences of trials generated by the models, but this data is not stored anywhere, and re-running the models (not all of which are publicly available anyway) seems excessive. Hence, we instead make mild assumptions to estimate conservative CIs around the empirical values, demonstrating the utility of our copy-model. (See Appendix˜I for details on this procedure.) We obtain the result depicted in Figure˜5 for the top-30 models (by EC to humans). Again, the best model is a CLIP-trained ViT. For the other models, the CIs are so large that any of the top ten models could be the second-best one, meaning that the true rank order between models remains unclear.
@@
Figure 5: Confidence Intervals for Brain-Score ECs. Like in Figure˜4, we plot the 95% confidence interval of the EC of the top 30 Brain-Score models. Note that the confidence intervals of many mid-range models overlap.
@@ 5 Discussion
In this work, we have built on the error consistency metric proposed by geirhos2020beyond, demonstrating how confidence intervals and significance tests for empirical measurements of EC can be calculated using bootstrapping techniques. Additionally, we have presented a new computational model of EC for planning sufficiently powerful, conclusive experiments. We have used our methodology to revisit influential Neuro-AI results schrimpf2018brainscore, finding that many results are not statistically reliable, because previous studies were underpowered. In the following, we recommend best practices for future benchmarking of behavioral alignment, and facilitate their implementation by making our python package publicly available. A core contribution of our work is an improved understanding of error consistency through a new computational model: EC is broken down into two interpretable components, copy probability $p_{copy}$ and marginal mismatch $f$. As an analogy, this is similar to breaking down accuracy into true positive and false positive rates—a necessary shift in perspective that will improve interpretations in future studies measuring similarities between classifiers.
5.1 Practical Recommendations: How to reliably benchmark behavioral alignment
@@
From our analysis, we can derive concrete recommendations for practitioners. As shown in Figure˜3, a sufficiently high number of trials is required to obtain small confidence intervals around EC values. As a rule of thumb, we suggest to collect at least 1000 trials per classifier to balance this requirement against practical constraints. In experiments with human observers, achieving very high resolution of EC differences could become prohibitively expensive, underlining the need for a good selection of stimuli, which probe differences efficiently. Since the size of CIs also depends on the accuracy mismatch between classifiers, we suggest to aim for an accuracy around 75%, but not more than 90% to avoid ceiling-effects, which can be severe. The same holds for floor-effects. This is in agreement with standard best practices from psychophysics, where one aims for a similar accuracy level in human observers to sustain motivation whilst avoiding ceiling effects. Finally, future work should always report confidence intervals and check results for statistical significance, which can easily be done with the python package we provide.
5.2 Conclusion
@@
The field of ML research relies heavily on the powerful machinery of benchmarking (hardt2025emerging). Hence, the correctness and reliability of benchmarks is crucial for progress, since ill-calibrated benchmarks will lead us astray. The field is also notorious for not quantifying uncertainty properly (pineau2021improving; miller2024adding; lehmann2025quantifying; bouthillier2021accounting), which is especially important in the context of benchmarks because of their large influence on other developments. Our work raises general questions about how benchmarks should aggregate results. For benchmarks that explicitly compute rankings, we give the concrete recommendations of (a) providing confidence intervals around individual estimates and (b) quantifying the stability of the resulting rankings as we did with error consistency.
@@
In this work, we have revisited results by geirhos2021partial which influence the Brain-Score benchmark schrimpf2018brainscore, revealing that many differences reported between deep vision models are not statistically reliable. The error consistency metric is particularly susceptible to such issues, because it is inherently noisy, so the requirements for stable measurements are hard to meet in practice. We envision that our work will help the field to overcome these issues, by providing tools to plan sufficiently powerful experiments.
@@
Funded, in part, by the Collaborative Research Centre (CRC) “Robust Vision – Inference Principles and Neural Mechanisms” of the German Research Foundation (DFG; SFB 1233), project number 276693517. FAW acknowledges funding by the BBVA Foundation Programme Grant “Harnessing Vision Science to Overcome the Critical Limitations of Artificial Neural Networks”. This work was additionally supported by the German Federal Ministry of Education and Research (BMBF): Tübingen AI Center, FKZ: 01IS18039A. WB acknowledges financial support via an Emmy Noether Grant funded by the German Research Foundation (DFG) under grant no. BR 6382/1-1 and via the Open Philanthropy Foundation funded by the Good Ventures Foundation. WB, FAW and KM are members of the Machine Learning Cluster of Excellence, EXC number 2064/1 – Project number 390727645. KM is supported by the Carl Zeiss Foundation through the CZS Institute for Artificial Intelligence and Law. This research utilized compute resources at the Tübingen Machine Learning Cloud, DFG FKZ INST 37/1057-1 FUGG. The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Thomas Klein.
We would like to thank Robert Geirhos for helpful discussions, and for providing us with his raw data for model-vs-human. We also thank all anonymous reviewers, who provided valuable feedback during the rebuttal period, which improved the final manuscript.
@@ Appendix A Related Work
Within the last decade, Deep Neural Networks (DNNs) have been proposed as models of the human visual system [doerig2023neuroconnectionist, kriegeskorte2015deep, cichy2019deep, kietzmann2017deep], not only because they form the only class of image-computable models that achieves human-level performance on benchmark tasks like ImageNet [russakovsky2015imagenet], but also because performance-optimized DNNs are the best predictors of biological neural activations [yamins2014performance, kubilius2019brain, zhuang2021unsupervised]. However, the suitability of DNNs as models of vision is debated [Bowers2022, wichmann2023deep, maniquet2024recurrent, serre2019deep], as is the question of how comparisons should best be conducted [firestone2020performance, lonnqvist2021comparative, wichmann2023deep].
@@
The underlying idea motivating error consistency is that of “molecular psychophysics” [green1963consistency], which is to go beyond aggregate measures like accuracy, and instead compare behavior on a trial-by-trial basis. EC promises to achieve this by building on Cohen’s $\kappa$ [cohen1960coefficient, Mart_n_Andr_s_2024], which considers the trial-level agreement observed for a pair of classifiers. geirhos2020beyond proposed EC in the context of human-machine comparison, but in principle, the method applies to the general setting of comparing arbitrary classifiers. Mart_n_Andr_s_2024 observe that $\kappa$ is a biased estimator and should be corrected, which we support in our implementation. safak2020min also observe the problem of $\kappa^{\prime}s$ dependency on the marginals, which we outline in Section˜2. They propose to scale $\kappa$ between its theoretical limits in an attempt to obtain a measure of error consistency that is orthogonal to accuracy. goel2025great propose to apply Cohen’s $\kappa$ directly to the responses given by both classifiers $\hat{y}_{i}$ rather than the correctness values $r_{i}$, and adapt the marginal probabilities of independent classifiers to reflect their accuracy. We still focus on the standard error consistency proposed by geirhos2020beyond because it has seen the widest application. In principle, our bootstrapping methodology also works for these other metrics.
@@
Benchmarking has a long history and is central to modern machine learning [koch2024protoscience, singh2025leaderboard, hardt2025emerging, orr2024ai]. For example, the popular Papers with Code database [^2] https://paperswithcode.com/sota lists approximately 14,000 different benchmarks. With the emergence of modern benchmarks and leaderboards, issues such as overfitting to benchmarks [singh2025leaderboard], the limitations of metrics [thomas2020problem], and various sources of uncertainty have been identified and challenged [lehmann2025quantifying, bouthillier2021accounting]. In general, it has been criticized that benchmarks should not be the primary goal of scientific machine learning research [alzahrani2024benchmarks]. At the same time, it has been pointed out that benchmarking itself requires a scientific approach and some kind of standard methodology [thiyagalingam2022scientific, sculley2018winner, hardt2025emerging]. Similar to our work, nado2021uncertainty argued that uncertainty is an important aspect to consider for benchmarks. In contrast to previous work, we focus on uncertainty in the behavioral comparison of classifiers using error consistency.
Appendix B Proof of Proposition˜2